# Spectra-to-Structure and Structure-to-Spectra Inference Across the Periodic Table

### v1

@@ Abstract
X-ray Absorption Spectroscopy (XAS) is a powerful technique for probing local atomic environments, yet its interpretation remains limited by the need for expert-driven analysis, computationally expensive simulations, and element-specific heuristics. Recent advances in machine learning have shown promise for accelerating XAS interpretation, but many existing models are narrowly focused on specific elements, edge types, or spectral regimes. In this work, we present XAStruct, a learning framework capable of both predicting XAS spectra from crystal structures and inferring local structural descriptors from XAS input. XAStruct is trained on a large-scale dataset spanning over 70 elements across the periodic table, enabling generalization to a wide variety of chemistries and bonding environments. The model includes the first machine learning approach for predicting neighbor atom types directly from XAS spectra, as well as a unified regression model for mean nearest-neighbor distance that requires no element-specific tuning. While we explored integrating the two pipelines into a single end-to-end model, empirical results showed performance degradation. As a result, the two tasks were trained independently to ensure optimal accuracy and task-specific performance. By combining deep neural networks for complex structure-property mappings with efficient baseline models for simpler tasks, XAStruct offers a scalable and extensible solution for data-driven XAS analysis and local structure inference. The source code will be released upon paper acceptance.
@@ 1 Introduction
To address these limitations, AI-driven approaches have emerged as promising alternatives to accelerate and automate the interpretation of XAS. Inspired by recent breakthroughs such as AlphaFold in protein structure prediction [9, 10, 11] and the broader success of AI in healthcare, astrophysics, and materials discovery [12, 13, 14, 15], these methods offer the potential for data-driven solutions that scale efficiently and adapt across chemical systems. Existing ML models for the prediction of XAS have focused mainly on crystalline inorganic materials [16, 17, 18, 19, 20], typically modeling either structure-to-spectra or spectra-to-structure mappings. However, many of these models are limited to a small number of elements, specific absorption edges, or narrow spectrum types, which restricts their generalization and practical utility.
Figure 1: Periodic Table Coverage in XAS2Struct. Green blocks indicate elements included in the task; white blocks are excluded due to missing or insufficient data.
This work presents XAStruct, a novel two-stage machine learning framework designed to enhance both the accuracy and completeness of XAS-related predictions. The first stage employs a physics principle-aware graph neural network (GNN) encoder (in this work, the CHGNet[21] backbone is utilized) that extracts complex, non-linear features from iron-containing crystal structures, enabling the accurate prediction of XANES and EXAFS spectra. The second stage goes beyond mere spectral prediction, it includes both classification and regression models to predict not only the type and number of neighboring atoms around the absorbing atom, such as Oxygen(O), Sulfur(S), and Copper(Cu), but also the mean nearest neighbor distance. As shown in Figure 1, XAStruct is trained on a large-scale dataset spanning over 70 elements across the periodic table, covering both K- and L-edge XANES and EXAFS regimes, which enables broad generalization to diverse chemistries.
@@
  • Enables learning in both structure-to-spectra and spectra-to-structure pipelines, within a unified framework trained across a wide range of absorbing atoms.
  • Achieves periodic table-wide generalization, enabled by training on over 70 elements, rather than focusing on a single element or material class.
@@
  • Offers a unified model for MNND prediction from spectra, serving as a robust surrogate for geometric interpretation and structure refinement.
@@ 2 Related Works
In contrast, our work introduces a GNN-based architecture built from scratch, designed for generalization across over 70 elements, and capable of predicting both K-edge and L-edge XANES and EXAFS spectra. By explicitly encoding elemental identity through masking and leveraging architectural components such as GatedLinear layers and SwiGLU activations, our model achieves strong generalization across edge types, energy regions, and diverse chemical environments. This approach marks a significant step beyond prior work focused on isolated systems, offering a scalable and interpretable framework for comprehensive XAS prediction.
@@
In contrast to the above work, ours introduces the first unified MNND prediction model capable of generalizing across more than 70 elements using shared weights. And to our knowledge, this work is the first to explore the prediction of neighbor atom types directly from XAS spectra. While CN and atom-type models are still trained per element due to categorical variance, our approach avoids reliance on handcrafted features and demonstrates that structural descriptors can be learned in a scalable, extensible manner, balancing interpretability, generalization, and precision in spectrum-based structure prediction.
@@ 3.1 X-ray Absorption Spectroscopy Overview
Figure 2: Illustration of X-ray Absorption Spectrum with XANES and EXAFS regions.
X-ray Absorption Spectroscopy (XAS) is a powerful technique for probing the local electronic and atomic structure of materials. Based on the Beer-Lambert law [28], the attenuation of X-ray intensity is governed by the linear absorption coefficient $\mu(E)$, which depends on photon energy $E$, sample thickness $d$, and material composition [29, 30]. For monatomic systems, an empirical relation approximates $\mu(E)\propto\rho Z^{4}/(AE^{3})$, where $Z$ and $A$ are atomic number and mass, respectively.
@@ 4 XAStruct: Bridging XAS and Structural Inference
(a) Structure-to-Spectrum: Predicting XANES/EXAFS from Crystal Graph
(b) Spectrum-to-Structure: Predicting CN/MNND/Neighbor Atom from XAFS
Figure 3: XAStruct: Bridging XAS and Structural Inference.
The analysis in the last section implies that traditional simulation or fitting-based techniques for predicting XAFS or material properties are computationally intensive and often limited in scalability, transferability, and automation. Furthermore, recovering local structural descriptors from spectra is particularly challenging due to the ill-defined and many-to-one nature of the problem. Although earlier machine learning models offered a data-driven alternative, they are often restricted to specific elements, edge types, or handcrafted features.
To address these challenges, we propose XAStruct—short for Bridging XAS and Structural Inference. Although the structure-to-spectrum and spectrum-to-structure tasks are trained independently, they represent two complementary directions of XAFS-driven learning in crystalline materials. As illustrated in Figure 3, XAStruct is composed of two distinct components: a forward pipeline that predicts XANES and EXAFS spectra from crystal graphs, and an inverse pipeline that infers structural descriptors directly from XAFS spectra. In the sections that follow, we describe each component in detail.
@@
  • GatedLinear: a linear projection with a gating branch, for input $\mathbf{x}\in\mathbb{R}^{d}$:
    $\mathtt{GatedLinear}(\mathbf{x})=(\mathbf{W}_{v}\mathbf{x}+\mathbf{b}_{v})\cdot\sigma(\mathbf{W}_{g}\mathbf{x}+\mathbf{b}_{g})$
@@
$\mathtt{SBlock}(\mathbf{x})=\mathtt{SwiGLU}(\mathtt{LayerNorm}(\mathtt{GatedLinear}(\mathbf{x})))$
And for a $k$-layer ($k-1$ blocks) SGMLP, it is constructed in this way:
$\mathtt{SGMLP}(\mathbf{x})=\mathtt{GatedLinear}\left(\mathtt{SBlock}^{(k-1)}\circ\cdots\circ\mathtt{SBlock}^{(1)}(\mathbf{x})\right)$ (6)
@@
$\mathtt{ConvBlock}(\mathbf{z}_{\mathbf{x},*})=\mathtt{AvgPooling}(\mathtt{ReLU}(\mathtt{BatchNorm}(\mathtt{Conv1D}(\mathbf{z}_{\mathbf{x},*}))))$ (11)
@@ 5 Experiment
While prior datasets [32, 33] have been used for XAS prediction, they do not contain sufficient input information (e.g., full 3D structure graphs or energy-resolved grids) required for evaluating our model. Therefore, we collected consistent structure-spectrum pairs across 70+ elements (see Figure 1 and Figure S10 for details), from Materials Project[34], which supports both forward and inverse learning tasks. Resulting in a dataset comprising over 43,000 unique crystal structures and approximately 120,000 corresponding X-ray absorption spectra (XANES and EXAFS), spanning both K and L edges across more than 70 chemical elements. The datasets are split into 8:2 subsets for training and validation. We conduct extensive experiments on the constructed dataset and train some models for each task under tailored settings to accommodate the complexity and variability inherent in different spectral and structural properties. The experiments are run on four NVIDIA RTX A6000-48GB GPUs.
@@ 5.1 Predicting K/L-edge XANES and EXAFS from Crystal Graphs
To account for the sensitivity of spectral features to the absorbing element and edge type, we train separate models for each element-edge-spectrum combination. This element-specific training avoids cross-element feature drift and allows models to specialize in chemically consistent environments. Models are trained using Mean Squared Error (MSELoss) between predicted and reference spectra, and optimized with AdamW (learning rate $10^{-4}$, weight decay 0.01).
@@ 5.2 Predicting Structural Descriptors from K-edge XAFS Spectra
Unified Model for MNND Prediction
For MNND prediction, we propose a single, unified SGMLP-Convolution hybrid regression model that accepts the full XAFS spectrum (including $\mu(E_{i})$ and $E_{i}$, as shown in Fig 3(b)) as input. The model is trained across all elements jointly, enabling it to generalize across more than 70 elements, including transition metals, main-group elements, and rare earths. Despite the chemical diversity in the training set, this model achieves robust performance without element-specific tuning, demonstrating remarkable generalizability when compared to other baseline models trained on element-specific datasets.
CN and Neighbor Atom Classification
In contrast to MNND, predicting CN and neighbor atom types raises a more discrete and element-specific classification challenge. CN values typically range over a small set (e.g., 2, 3, 4, 5, 6, 8, 12), but vary significantly across different elements and crystal classes. Neighbor atom classification is even more difficult due to the high variance (over 70 element types) and sparsity of classes in any given crystal structure. For example, while there are 70+ possible atom types in the full dataset, any specific XAFS spectrum is typically associated with only 2-4 neighboring species, resulting in highly imbalanced and underrepresented labels.
Baseline Justification and Training Setup
As shown in Table 2 and Figures S8, S9, and S6 (supplementary), XAStruct achieves strong performance across all structure-related prediction tasks. The unified MNND model generalizes well across elements and bond distances, capturing the underlying geometric trends with high fidelity. In contrast, as illustrated in Figure S7, the random forest baseline exhibits systematic errors—overestimating MNND for tightly packed structures and underestimating it for loosely coordinated environments, resulting in a lower regression slope (0.92) compared to XAStruct (0.97). For neighbor atom classification, XAStruct achieves notably higher precision and recall, especially in rare or chemically diverse cases, highlighting its advantage over conventional models. While the random forest outperforms MLP in CN prediction, its lightweight nature also yields faster training and inference, making it a practical baseline. Overall, these results validate our design choices and emphasize the benefits of combining task-specific architectures with robust learning objectives.
@@
5.3 Exploring Pipeline Integration
We explored whether coupling the structure-to-spectrum and spectrum-to-structure pipelines could improve performance through two strategies. First, we added a cosine similarity loss between the structural embedding $\mathbf{z}_{\text{struct}}$ (from the forward model) and the spectral embedding $\mathbf{z}_{\mathbf{x},*}$ (from the inverse model) to encourage representational alignment. Second, we injected $\mathbf{z}_{\text{struct}}$ into the inverse model either by concatenation or as a substitute for learned spectral features.
While these strategies were conceptually attractive, they led to limited or even negative performance impact. MNND prediction showed marginal improvements (e.g., $\sim$0.002 lower MAE for certain elements), but CN and neighbor atom classification degraded by roughly 2% in accuracy and F1 score. We attribute this to representational mismatch: structural and spectral embeddings encode fundamentally different distributions, and alignment introduces regularization that hinders task-specific specialization. Additionally, directly injecting structure-derived embeddings likely biases the inverse model toward geometries it cannot fully infer from spectra alone.
These results suggest that the two tasks require distinct representational pathways and optimization goals. Thus, training the pipelines independently remains the most robust strategy for generalizing across the wide chemical and spectral diversity of XAS data.
@@ 6 Limitations and Conclusion
However, several limitations remain. While XAStruct supports dual learning objectives, the two pipelines are trained independently. Our experiments to connect them, either by aligning embeddings or using one pipeline’s output as input to the other, led to degraded performance, likely due to differing representational requirements and noise sensitivity between the tasks. Additionally, CN and neighbor atom prediction models must be trained separately per element, as unified models struggle with class imbalance and high categorical variance. Our structural predictions still require expert interpretation to reconstruct full geometries, limiting full automation. Moreover, although the MNND model generalizes well, CN and neighbor atom predictions are bounded by closed-set assumptions and limited resolution in fine spectral features such as XANES pre-edges and EXAFS oscillations. These challenges highlight promising directions for future work, including cross-modal alignment, few-shot generalization for underrepresented chemistries, and physically grounded model regularization to enhance robustness and interpretability.
Appendix A Technical Appendices and Supplementary Material

### v2

@@ Abstract
X-ray Absorption Spectroscopy (XAS) is a powerful technique for probing local atomic environments, yet its interpretation remains limited by the need for expert-driven analysis, computationally expensive simulations, and element-specific heuristics. Recent advances in machine learning have shown promise for accelerating XAS interpretation, but many existing models are narrowly focused on specific elements, edge types, or spectral regimes. In this work, we present XAStruct, a learning-based system capable of both predicting XAS spectra from crystal structures and inferring local structural descriptors from XAS input. XAStruct is trained on a large-scale dataset spanning over 70 elements across the periodic table, enabling generalization to a wide variety of chemistries and bonding environments. The framework includes the first machine learning approach for predicting neighbor atom types directly from XAS spectra, as well as a generalizable regression model for mean nearest-neighbor distance that requires no element-specific tuning. By combining deep neural networks for complex structure–property mappings with efficient baseline models for simpler tasks, XAStruct offers a scalable and extensible solution for data-driven XAS analysis and local structure inference. The source code will be released upon paper acceptance.
@@ 1 Introduction
To address these limitations, AI-driven approaches have emerged as promising alternatives to accelerate and automate the interpretation of XAS. Inspired by recent breakthroughs such as AlphaFold in protein structure prediction [9, 10, 11] and the broader success of AI in healthcare, astrophysics, and materials discovery [12, 13, 14, 15, wang2025x], these methods offer the potential for data-driven solutions that scale efficiently and adapt across chemical systems. Existing ML models for the prediction of XAS have focused mainly on crystalline inorganic materials [16, 17, 18, 19, 20], typically modeling either structure-to-spectra or spectra-to-structure mappings. However, many of these models are limited to a small number of elements, specific absorption edges, or narrow spectrum types, which restricts their generalization and practical utility.
This work presents XAStruct, a novel two-stage machine learning framework designed to enhance both the accuracy and completeness of XAS-related predictions. The first stage employs a physics principle-aware graph neural network (GNN) encoder (in this work, the CHGNet [21] backbone is utilized) that extracts complex, non-linear features from crystal structures, enabling the accurate prediction of XANES and EXAFS spectra. The second stage goes beyond mere spectral prediction, it includes both classification and regression models to predict not only the type and number of neighboring atoms around the absorbing atom, such as Oxygen (O), Sulfur (S), and Copper (Cu), but also the mean nearest neighbor distance. As shown in Supplementary Figure S12, XAStruct is trained on a large-scale dataset spanning over 70 elements across the periodic table, covering both K- and L-edge XANES and EXAFS regimes, which enables broad generalization to diverse chemistries.
@@
  • Enables learning in both structure-to-spectra and spectra-to-structure pipelines, within a comprehensive framework trained across a wide range of absorbing atoms.
  • Achieves periodic table–wide generalization, enabled by training on over 70 elements, rather than focusing on a single element or material class.
@@
  • Offers a generalizable model for MNND prediction from spectra, serving as a robust surrogate for geometric interpretation and structure refinement.
@@ 2 Related Works
In contrast, our work introduces a learning framework built around a GNN-based architecture designed for generalization across over 70 elements, and capable of predicting both K-edge and L-edge XANES and EXAFS spectra. By explicitly encoding elemental identity through masking and leveraging architectural components such as GatedLinear layers and SwiGLU activations, our model achieves strong generalization across edge types, energy regions, and diverse chemical environments. This approach marks a significant step beyond prior work focused on isolated systems, offering a scalable and interpretable framework for comprehensive XAS prediction.
@@
In contrast to the above work, ours introduces the first generalizable MNND prediction model capable of generalizing across more than 70 elements using shared weights. And to our knowledge, this work is the first to explore the prediction of neighbor atom types directly from XAS spectra. While CN and atom-type models are still trained per element due to categorical variance, our approach avoids reliance on handcrafted features and demonstrates that structural descriptors can be learned in a scalable, extensible manner, balancing interpretability, generalization, and precision in spectrum-based structure prediction.
@@ 3.1 X-ray Absorption Spectroscopy Overview
Figure 1: X-ray absorption spectrum with XANES (near-edge) and EXAFS (oscillatory) regions.
X-ray Absorption Spectroscopy (XAS) is a powerful technique for probing the local electronic and atomic structure of materials. Based on the Beer–Lambert law [28], the attenuation of X-ray intensity is governed by the linear absorption coefficient $\mu(E)$, which depends on photon energy $E$, sample thickness $d$, and material composition [29, 30]. For monatomic systems, an empirical relation approximates $\mu(E)\propto\rho Z^{4}/(AE^{3})$, where $Z$ and $A$ are atomic number and mass, respectively.
@@ 4 XAStruct: Bridging XAS and Structural Inference
Figure 2: Structure to Spectrum: Predicting XANES/EXAFS from Crystal Graph
Figure 3: Spectrum to Structure: Predicting CN/MNND/Neighbor Atom from XAFS
While powerful, traditional XAS analysis is often limited by computationally intensive simulations and the ill-posed nature of inverting spectra to find structure. Existing machine learning models often lack generalizability, being restricted to specific elements or spectral regions.
To address these challenges, we propose XAStruct—short for Bridging XAS and Structural Inference. Although the structure-to-spectrum and spectrum-to-structure tasks are trained independently, they represent two complementary directions of XAFS-driven learning in crystalline materials. As illustrated in Figure 2,3, XAStruct is composed of two distinct components: a forward pipeline that predicts XANES and EXAFS spectra from crystal graphs, and an inverse pipeline that infers structural descriptors directly from XAFS spectra. In the sections that follow, we describe each component in detail.
@@
  • GatedLinear(GL): a linear projection with a gating branch, for input $\mathbf{x}\in\mathbb{R}^{d}$:
    $\mathtt{GL}(\mathbf{x})=(\mathbf{W}_{v}\mathbf{x}+\mathbf{b}_{v})\cdot\sigma(\mathbf{W}_{g}\mathbf{x}+\mathbf{b}_{g})$
@@
$\mathtt{SBlock}(\mathbf{x})=\mathtt{SwiGLU}(\mathtt{LayerNorm}(\mathtt{GL}(\mathbf{x})))$
And for a $k$-layer (k1k-1 blocks) SGMLP, it is constructed in this way:
$\mathtt{SGMLP}(\mathbf{x})=\mathtt{GL}\left(\mathtt{SBlock}^{(k-1)}\circ\cdots\circ\mathtt{SBlock}^{(1)}(\mathbf{x})\right)$ (6)
@@
$f_{\mathrm{conv}}(\mathbf{z}_{\mathbf{x},*})=\mathtt{AvgPool}(\mathtt{ReLU}(\mathtt{BN}(\mathtt{Conv1D}(\mathbf{z}_{\mathbf{x},*}))))$ (11)
@@ 5 Experiment
While prior datasets [32, 33] have been used for XAS prediction, they do not contain sufficient input information (e.g., full 3D structure graphs or energy-resolved grids) required for evaluating our model. Therefore, we collected consistent structure-spectrum pairs across 70+ elements (see Figure LABEL:fig:pt_main and Figure S12 for details), from Materials Project[34], which supports both forward and inverse learning tasks. Resulting in a dataset comprising over 43,000 unique crystal structures and approximately 120,000 corresponding X-ray absorption spectra (XANES and EXAFS), spanning both K and L edges across more than 70 chemical elements. The datasets are split into 8:2 subsets for training and validation. We conduct extensive experiments on the constructed dataset and train some models for each task under tailored settings to accommodate the complexity and variability inherent in different spectral and structural properties. The experiments are run on four NVIDIA RTX A6000-48GB GPUs.
@@ 5.1 Predicting K/L-edge XANES and EXAFS from Crystal Graphs
To account for the sensitivity of spectral features to the absorbing element and edge type, we train separate models for each element–edge–spectrum combination. This element-specific training avoids cross-element feature drift and allows models to specialize in chemically consistent environments. Models are trained using Mean Squared Error (MSELoss) between predicted and reference spectra, and optimized with AdamW (learning rate $10^{-4}$, weight decay 0.01).
@@ 5.2 Predicting Structural Descriptors from K-edge XAFS Spectra
Table 3: Ablation study of XAStruct vs. variants with key components removed. Averages are over the validation set.
Structure $\to$ Spectrum Spectrum $\to$ Structure
Model Configuration XANES (K) MAE MNND MAE Neighbor Atom F1 (%) Neighbor Atom Acc. (%)
XAStruct (Ours) $\textbf{0.0537}_{\pm 0.0005}$ $\textbf{0.0350}_{\pm 0.0007}$ $\textbf{88.76}_{\pm 0.01}$ $\textbf{92.96}_{\pm 0.02}$
SwiGLU (use ReLU) $0612_{\pm 0.0006} 0.0$$0421_{\pm 0.0005}$ $85.15_{\pm 0.01}$ $90.12_{\pm 0.01}$
GatedLinear (use nn.Linear) $0595_{\pm 0.0007} 0.0$$0403_{\pm 0.0005}$ $86.32_{\pm 0.01}$ $91.05_{\pm 0.02}$
Standard MLP $0689_{\pm 0.0008} 0.0$$0498_{\pm 0.0007}$ $3.45_{\pm 0.01} 88$$8.54_{\pm 0.03}$
Generalizable Model for MNND Prediction
For MNND prediction, we propose a single, generalizable SGMLP-Convolution hybrid regression model that accepts the full XAFS spectrum (including $\mu(E_{i})$ and $E_{i}$, as shown in Fig 3) as input. The model is trained across all elements jointly, enabling it to generalize across more than 70 elements, including transition metals, main-group elements, and rare earths. Despite the chemical diversity in the training set, this model achieves robust performance without element-specific tuning, demonstrating remarkable generalizability when compared to other baseline models trained on element-specific datasets.
CN and Neighbor Atom Classification
In contrast to MNND, predicting CN and neighbor atom types raises a more discrete and element-specific classification challenge. CN values typically range over a small set (e.g., 2, 3, 4, 5, 6, 8, 12), but vary significantly across different elements and crystal classes. Neighbor atom classification is even more difficult due to the high variance (over 70 element types) and sparsity of classes in any given crystal structure. For example, while there are 70+ possible atom types in the full dataset, any specific XAFS spectrum is typically associated with only 2–4 neighboring species, resulting in highly imbalanced and underrepresented labels.
Baseline Justification and Training Setup
As shown in Table 2 and Figures S8, S10, and S6 (supplementary), XAStruct achieves strong performance across all structure-related prediction tasks. The generalizable MNND model generalizes well across elements and bond distances, capturing the underlying geometric trends with high fidelity. In contrast, as illustrated in Figure S7, the random forest baseline exhibits systematic errors—overestimating MNND for tightly packed structures and underestimating it for loosely coordinated environments, resulting in a lower regression slope (0.92) compared to XAStruct (0.97). For neighbor atom classification, XAStruct achieves notably higher precision and recall, especially in rare or chemically diverse cases, highlighting its advantage over conventional models. While the random forest outperforms MLP in CN prediction, its lightweight nature also yields faster training and inference, making it a practical baseline. Overall, these results validate our design choices and emphasize the benefits of combining task-specific architectures with robust learning objectives.
5.3 Ablation Study on Architectural Components
To validate the design choices of our architecture, we conducted an ablation study focusing on the key components of our SGMLP module, which is central to both the forward and inverse pipelines. The study aims to quantify the performance impact of the GatedLinear layer and the SwiGLU activation function. We evaluated four model configurations on the most representative tasks: K-edge XANES prediction, MNND prediction, and neighbor atom classification.
The configurations are as follows:
  • XAStruct (Ours): The full proposed model utilizing the complete SGMLP block (GatedLinear layers with SwiGLU activation).
  • SwiGLU: A variant where the SwiGLU activation in the SGMLP is replaced with a standard ReLU activation, while keeping the GatedLinear layer.
  • GatedLinear: A variant where the GatedLinear layer is replaced with a standard nn.Linear layer, while keeping the SwiGLU activation.
  • Standard MLP: A baseline model where the entire SGMLP block is replaced by a conventional Multi-Layer Perceptron composed of standard nn.Linear layers and ReLU activation functions.
The results, summarized in Table 3, clearly demonstrate the benefits of our chosen components. In all evaluated tasks, the full XAStruct model achieves the best performance. Removing either the SwiGLU activation or the GatedLinear layer leads to a noticeable degradation in performance. Replacing the GatedLinear layer with a standard linear layer while keeping SwiGLU (- GatedLinear) results in a moderate drop in accuracy and F1-score, suggesting that the gating mechanism is effective at controlling information flow.
More significantly, replacing SwiGLU with ReLU causes a more substantial performance decrease across all metrics. This highlights the advantage of SwiGLU’s non-linear, data-dependent activation for capturing the complex relationships in spectral data. Finally, the standard MLP baseline shows the poorest performance, confirming that the combination of both gating and the advanced activation function in our SGMLP architecture is crucial for achieving state-of-the-art results.
@@ 6 Limitations and Conclusion
However, several limitations remain. While XAStruct supports dual learning objectives, the two pipelines are trained independently. Additionally, CN and neighbor atom prediction models must be trained separately per element, as unified models struggle with class imbalance and high categorical variance. Our structural predictions still require expert interpretation to reconstruct full geometries, limiting full automation. Moreover, although the MNND model generalizes well, CN and neighbor atom predictions are bounded by closed-set assumptions and limited resolution in fine spectral features such as XANES pre-edges and EXAFS oscillations. These challenges highlight promising directions for future work, including cross-modal alignment, few-shot generalization for underrepresented chemistries, and physically grounded model regularization to enhance robustness and interpretability.
Appendix A Appendix